# A multi-task deep learning approach for lane-level pavement performance prediction with segment-level data

### v1

@@ 3.2.2 Multiple task-specific heads
The multiple task-specific heads, module (b) in Figure 2, is designed to generate predictions for each lane, allowing the model to learn unique pavement performance decay patterns. Each head comprises an LSTM unit and a fully connected layer that employs the Leaky Rectified Linear Unit (LeakyReLU) activation function. The LSTM component is instrumental in predicting pavement performance by analyzing the time series data derived from the LSTM-based shared layer. The fully connected layer that follows serves a dual purpose: it maps the LSTM’s predictions to a tailored range or space, and enhances the model’s capacity to distill and articulate a rich tapestry of features.
$y_{k}=g\left(\sum_{j=1}^{J}w_{jk}x_{j}+b_{k}\right)$ (8)
Where $g$ is activation function, $J$ is the number of neurons, $w_{jk}$ is the weight from neuron $j$ of the previous layer to neuron $k$, $x_{j}$ is the input from neuron $j$, and $b_{k}$ is the bias for neuron $k$.
In deep neural networks, the gradients may gradually decrease during back propagation, making the model difficult to train. The LeakyReLU activation function has a non-zero gradient for negative inputs, which helps the gradients propagate more effectively in the network, thereby accelerating the training and convergence of the model. The mathematical expression for the LeakyReLU is:
$\hfill F(x)=\begin{cases}x,&\text{if }x\geq 0\\ \alpha x,&\text{if }x<0\end{cases}\hfill$ (9)
Where $\alpha$ is a small constant, often chosen between 0.01 and 0.1.
@@ 3.2.3 Multiple task-specific output layers
Multiple task-specific output layers, module (c) in figure 2, are used to transform the concatenation result, a two-dimensional tensor consisting of initial prediction and auxiliary features of lanes, into final lane-level predictions. Each task-specific output layer is responsible for a specific lane.
$\mathbf{\widehat{Y}_{n}}=\mathbf{X}\times\mathbf{W_{n}}$ (10)
$\mathbf{\widehat{Y}_{n}}=\left[\widehat{\boldsymbol{y}}_{n1},\widehat{\boldsymbol{y}}_{n2},\ldots,\widehat{\boldsymbol{y}}_{n\rho}\right]$ (11)
Where $\mathbf{\widehat{Y}_{n}}$ is the final predictions of lane $n$, $n$ is the lane number, $\mathbf{X}$ is the concatenation result, $\mathbf{W_{n}}$ is the weight matrix of task-specific output layer for lane $n$, and $\rho$ is the number of prediction units.
@@ 3.2.4 Total loss calculation
In a multi-task learning-based model, the loss function feedback mechanism balances the training of various tasks, ensuring that each task benefits from improvements. In our model, the loss function is the aggregate of the losses from each pavement prediction task. Once the total loss is calculated, parameter updates are executed through the gradients derived from the loss function.
$L_{Total}=L_{1}+L_{2}+\dots+L_{N}$ (12)
$L_{n}=\frac{1}{\rho}\sum_{i=1}^{\rho}(y_{i}-\hat{y}_{i})^{2}$ (13)
Where $L_{Total}$ is the total loss of model training, $N$ is the number of lanes, $L_{n}$ is the MSE loss for the pavement prediction of lane n, $y_{i}$ is the actual value of prediction unit $i$, and $\hat{y}_{i}$ is the predicted value of prediction unit $i$.
4 REULSTS
@@ 4.1 Dataset and preprocessing
The pavement inspection dataset from Zhengzhou, Jiaozuo, and Luoyang in Henan Province, China, covering 2020 to 2023, was used for model verification. The data from 2020 to 2022 is segment-level and recorded in units of 1000m (around 0.6 mile); the data from 2023 is lane-level and recorded in units of 100m (around 0.06 mile). The dataset records road names, the direction of travel (upstream and downstream), milepost numbers, pavement performance indicators (such as PCI, PQI, and RQI), pavement surface material, etc. A comprehensive introduction of features is delineated in TABLE 1. To verify the general effectiveness of the model in different task quantities and indicator predictions, we establish three distinct datasets, 2-lane, 3-lane, and 4-lane, each dataset consisting of the road with a corresponding number of lanes. Each dataset will be used to predict PCI, PQI, and RQI.
Figure 3: The description of study case in Henan province, China
@@ 3 RQI Riding Quality Index Continuous From 0 to 100
5 Dir. Upstream or downstream direction of road Discrete Upstream, downstream
@@ 11 PSM pavement surface material Discrete Asphalt, concrete
12 LN Lane number, The innermost is lane 1 Discrete From 1 to 4
@@
All available datasets are collected from 2020 to 2023 and matched based on road name, direction, and mileposts. Consequently, time series can be formed for three pavement performance indicators, including PCI, PQI, and RQI. Only the time series that demonstrate a year-by-year degradation in pavement performance are retained. Because when pavement is not maintained, its performance is expected to continue declining. We do not take into account performance predictions in the presence of road maintenance disturbances. The final number of time series is presented in TABLE 2. Each dataset, including time series for PCI, PQI, and RQI and auxiliary features, will be partitioned into training and testing subsets with a ratio of 70:30. Features excluding PCI, PQI, and RQI in each training set will undergo independent Z-score normalization. And the same features in the corresponding test set will then be standardized using the mean and variance derived from their respective training sets. This approach ensures that the model training phase does not inadvertently incorporate test set information and guarantees that the scale and distribution of the test data match those of the training data. This alignment facilitates a fair and accurate assessment of the model’s performance.
Table 2: The data samples in three road scenarios
@@
4.2 Evaluation metrics
Mean absolute percentage error(MAPE) is employed to assess the performance of our model. MAPE reflects the average relative deviation of the predicted values from the actual values and is not influenced by the magnitude of the data. The smaller their values, the better the performance of the model. The calculation formula is as follows:
$MAPE_{n}=\frac{1}{m}\sum_{j=1}^{m}\left|\frac{y_{nj}-\hat{y}_{nj}}{y_{nj}}\right|\times 100\%$ (14)
$MAPE_{Total}=\frac{1}{N}\sum_{n=1}^{N}MAPE_{n}$ (15)
Where $n$ is the lane number, $MAPE_{n}$ is the MAPE of lane $n$, $m$ is the number of prediction units in lane $n$, $\hat{y}_{nj}$ is the ${j}^{th}$ unit’s predicted value of lane $n$, $y_{nj}$ is the ${j}^{th}$ unit’s actual value of lane $n$, $MAPE_{Total}$ is the MAPE of all lanes, and $N$ is the number of lanes.
@@ 4.3 Model setting
The model parameters can be set as follows. The LSTM-based shared layer, module(a) in Figure 2, includes two layers with a hidden layer of dimension 128. In the multiple task-specific heads, module(b) in Figure 2, the LSTM contains two layers with a hidden layer of dimension 64, and a fully connected layer has two layers with a hidden layer of dimension 64 and an output layer of dimension 32. The activation function employed in the fully connected layers is the LeakyReLU. The number of task-specific heads is variable, aligning with the number of lanes, and ranges from two to four. The fully connected layer in the multiple task-specific output layers is of size 32 and output dimension 1. The hyperparameters that underpin our model include, but are not limited to, the learning rate, batch size, and the number of training epochs. In our experiment, the learning rate is 0.001, the batch size is 32, and the model is subjected to 200 training epochs. The output of each layer is shown in Table 3.
@@ 4.4 Model performance
The following models are selected to compare the performance of our model:
(1) LSTM. It can effectively handle the long-term dependencies in sequential data.
(2) GRU. Gated Recurrent Unit. It is similar to LSTM, which also introduces a gating mechanism to control the flow of information. Compared with LSTM, it has a simpler structure, fewer parameters, and higher computational efficiency.
(3) XGBoost Regressor. eXtreme Gradient Boosting Regressor. It is a robust gradient-boosting decision tree algorithm. When making predictions, each historical value of the performance indicators is added as a feature.
(4) RF. Random Forest is an ensemble learning algorithm that builds multiple decision trees and combines their results to make the final prediction. When making predictions, each historical value of the performance indicators is added as a feature.
Two distinct experimental configurations were established, referred to as the ’lane-specific model’ and the ’mix model’, respectively, based on the four previously mentioned models. The lane-specific model will train a dedicated model for each lane. For instance, this strategy culminates four models within the context of a 4-lane road dataset. So, the training process of each model is not influenced by the decay patterns of other lanes. The mix model will add lane number as a feature, which is used to distinguish the decay patterns of different lanes. The model enables the concurrent generation of all lanes’ predictive outputs.
The proposed model and comparative models were deployed to predict lane-level PCI, PQI, and RQI using a 2-lane, 3-lane, and 4-lane dataset. Our model demonstrates superior performance when evaluated under the MAPE metric across all three datasets and for each performance indicator: PCI, PQI, and RQI. This out-performance underscores the efficacy of our model in delivering more precise lane-level predictions. The detailed MAPE results are presented in Table 4.
@@
(1) Our model exhibited the most optimal performance on the 4-lane dataset, trailed by the 3-lane dataset, while the 2-lane dataset manifested the poorest performance. The MAPE was 4.936%, 5.816%, and 10.309% respectively. This phenomenon is associated with the quantity of prediction units and tasks. The greater the number of units, the more information the model acquires. Furthermore, a greater number of lanes is more conducive to enhancing the shared information.
(2) Compared with other models, our model has accomplished an average absolute maximum reduction in MAPE of 3.092%, 2.699%, and 2.774% for the 4-lane, 3-lane, and 2-lane datasets, respectively. Correspondingly, the average absolute minimum reduction in MAPE is 1.480%, 0.556%, and 0.560% for the datasets mentioned above in sequence. The outcomes manifest the effectiveness of our model in leveraging related tasks to enhance the overall forecasting accuracy in pavement performance analysis.
(3) Our model attains an average MAPE of 7.852%, 7.275%, and 5.934%, respectively, for predicting PQI, PCI, and RQI. Furthermore, as depicted in Figure 1, the dataset exhibits the maximum average PQI difference of 2.834 between different lanes, succeeded by PCI with a difference of 1.950 and RQI with a difference of 0.946. The MAPE demonstrates a specific positive correlation with the performance disparity between lanes. The positive correlation indicates that the similarity between lanes and the resemblance between tasks can mitigate the model’s prediction error.
(4) Figure 4 demonstrates lane-level MAPE results using our model. Moreover, as the number of lanes escalates, the difference in MAPE among lanes narrows. Additionally, the MAPE of lane 1 and lane 2 generally shows a downward trend because the information of the newly added prediction tasks will improve the accuracy of the existing performance prediction tasks.
@@
(5) FIGURE 5, 6, and 7 show lane-level MAPE based on 2-lane, 3-lane, and 4-lane datasets. In most cases, the MAPE of each lane is the lowest using our model compared with the lane-specific and mix models. In the 2-lane dataset, there are 3 times, accounting for 50%, that the lane-level MAPE is not the lowest. And all are from lane 2. But in 3-lane and 4-lane datasets, the numbers that the lane-level MAPE is not the lowest are 2 and 4, only accounting for 22% and 33%.
@@ 4.5 Ablation experiment
An ablation experiment on the model structure and auxiliary features was conducted to assess the distinct contributions of each component on our 4-lane dataset. Regarding the model structure, three different test models were implemented. As for the auxiliary features, each feature was removed individually to assess its impact.
There are test models’ introductions. In test model 1, the LSTM-based shared layer is removed, so the pavement performance time series are sent directly to each task-specific head. The shared layer typically extracts common features from the input data, offering a universally applicable prediction outcome for all lanes. Furthermore, these layers incorporate a regularization effect that helps prevent over-fitting to any single task. In test model 2, the model eliminates the multiple task-specific heads. Instead, the output from the LSTM-based shared layer is transformed into a 2-dimensional format and then horizontally concatenated with auxiliary features. The presence of task-specific heads, each focusing on the prediction of a specific lane, can enhance predictive performance by concentrating on the relevant features and patterns most indicative of the outcomes for each lane. This specialization improves the accuracy and reliability of the predictions. In test model 3, the horizontal concatenation process is removed. The original model integrates auxiliary features that enhance its ability to represent input data. This integration is conducive to capturing the varied impacts on pavement performance deterioration more comprehensively, including factors such as traffic load and pavement material. It enables the model to detect more subtle and complex patterns. Moreover, feature integration provides increased regularization, which helps to mitigate the risk of over-fitting.
Our model and test models are used to predict PCI, PQI, and RQI, and the MAPE results are shown in Figure 8(a). The results indicate that all the components contribute positively to enhancing the accuracy of the predictions. The LSTM-based shared layer is of the most significant importance to our model. When this shared layer is removed, the MAPE increases by 1.631%, 1.611%, and 1.806% when predicting PCI, PQI, and RQI, respectively. To validate the impact of every auxiliary feature on the model’s predictive performance, each auxiliary feature was removed individually, and the MAPE results are shown in Figure 8(b). The results indicate that all the utilized auxiliary features contribute positively to enhancing the accuracy of the predictions. Also, there is not one feature that will greatly improve prediction accuracy. The reduction of MAPE ranges from 0.227% to 0.547% when predicting PCI, 0.300% to 0.697% when predicting PQI, and 0.245% to 0.663%.

### v2

@@ 3.2.2 Multiple task-specific heads
The module (b) in Figure 2 provides multiple task-specific heads, designed to learn unique pavement performance deterioration patterns for each lane. Each head comprises an LSTM unit and a fully connected neural network layer. The LSTM component is important in capturing pavement performance of each lane by taking hidden output and cell states from the LSTM-based shared layer. The following fully connected neural network layer maps the LSTM predictions to a tailored shape.
$\mathbf{S_{n}}=\mathbf{s_{n}}*\mathbf{W_{x}},\mathbf{S_{n}\in\mathbb{R}^{M\times V}},\mathbf{s_{n}\in\mathbb{R}^{M\times 1}}$ (8)
where,
$\mathbf{{S}_{n}}$ is the fully connected layer’s output at lane $n$;
$\mathbf{{s}_{n}}$ is the LSTM output at lane $n$;
$\mathbf{W_{s}}$ is the weight matrix of the fully connected neural network layer;
$V$ is the shape size of one output dimension in the fully connected neural network layer.
In the fully connected neural network layer, the LeakyReLU as activation function is used, as in equation 9. Since such activation function has a non-zero gradient for negative inputs, which makes the gradients propagation more effectively in the deep neural network.
$\hfill LReLU(Val)=\begin{cases}Val,&\text{if }Val\geq 0\\ \alpha Val,&\text{if }Val<0\end{cases}\hfill$ (9)
where,
$\alpha$ is a small constant, often ranging from 0.01 and 0.1.
@@ 3.2.3 Multiple task-specific output layers
The module (c) in Figure 2 first horizontally concatenate the output of module (b) and auxiliary features. Then the fully connected neural network layer transforms the concatenation tensor into lane-level predictions.
$\mathbf{X}=\left[\mathbf{X_{1}},\mathbf{X_{2}},\ldots,\mathbf{X_{m}},\ldots,\mathbf{X_{M}}\right],~{}\mathbf{X_{m}}\in\mathbf{R}^{U\times 1},~{}\mathbf{X}\in\mathbb{R}^{M\times U}$ (10)
$\mathbf{S}=\left[\mathbf{S_{1}},\ldots,\mathbf{S_{n}},\ldots,\mathbf{S_{N}},\mathbf{X}\right],~{}\mathbf{S_{n}}\in\mathbf{R}^{M\times V}$ (11)
$\mathbf{\widehat{Y}^{t+1}_{n}}=\mathbf{S}\times\mathbf{W_{n}},~{}\mathbf{\widehat{Y}^{t+1}_{n}}\in\mathbf{R}^{M\times 1}$ (12)
where,
$\mathbf{X}$ is the matrix of auxiliary features measured at all $M$ segments;
$U$ is the number of auxiliary features;
$\mathbf{S}$ is the concatenation result of module (b) output and auxiliary features;
$\mathbf{\widehat{Y}^{t+1}_{n}}$ is the final pavement performance predictions of lane $n$ at time step $t+1$;
$\mathbf{W_{n}}$ is the weight matrix in the fully connected neural network layer for lane $n$.
@@ 3.2.4 Total loss calculation
In the proposed multi-task learning-based model, the parameter estimation relies on the training loss of not only each lane-specific prediction, but also the whole sample prediction. In this study, mean squared error (MSE) is used as the loss function for each lane-specific prediction, to derive a total summation of all MSEs as the total loss.
$L=L_{1}+L_{2}+\dots+L_{n}+\dots+L_{N}$ (13)
$L_{n}=\frac{1}{M}\sum_{j=1}^{M}(\boldsymbol{y}^{t+1}_{n,~{}j}-\widehat{\boldsymbol{y}}^{t+1}_{n,~{}j})^{2}$ (14)
where,
$L$ is the overall pavement prediction MSE for model training;
$L_{n}$ is the pavement prediction MSE of lane n;
$\boldsymbol{y}^{t+1}_{n,~{}j}$ is the actual pavement performance on spatial prediction unit $j$ of lane $n$ at time step $t+1$.
4 Results
@@ 4.1 Dataset and preprocessing
One real pavement performance measurement dataset is collected in three cities of Zhengzhou, Jiaozuo, and Luoyang in Henan Province, China, as shown in Figure 3. The pavement performance is mainly evaluated with three common indicators including PCI, PQI, and RQI, ranging from the year of 2020 to the year of 2023. The data in the first three years is measured on each road segment with a length of around $1000m$ or 0.6 mile. The data in the last year is measured on each lane, dividing into units with a length of $100m$ or 0.06 mile. Meanwhile, a subset of auxiliary features has been collated, including road names, direction(upstream or downstream), road classification, city and town information, pavement surface material, traffic volume, and speed limit. All features, as well as their statistics, are summarized in TABLE 1.
Figure 3: The description of the study case in Henan Province, China
@@ 3 RQI Riding Quality Index Continuous From 0 to 100
5 Dir. Upstream or downstream direction of the road Discrete Upstream, downstream
@@ 11 PSM pavement surface material Discrete Asphalt, concrete
12 LN Lane number, and the innermost is lane 1 Discrete From 1 to 4
@@
Due to the data quality and maintenance, data cleaning was applied to the real dataset. First, deleting the road segments with the implementation of maintenance during data collection periods. Since the maintenance will significantly change the pavement performance in a short time. Second, measurement errors may result in fluctuations, violating year-by-year degradation pattern. The road segments with such fluctuations are also excluded from samples. In final, the whole dataset is divided into three distinct datasets based on three different road scenarios of one-way 2-lane, 3-lane, and 4-lane. The statistics of all three datasets are summarized in TABLE 2. Each dataset will be partitioned into training and testing set by a ratio of 7:3. The auxiliary features will undergo Z-score normalization before model training and testing.
Table 2: The number of data samples in three road scenarios
@@
4.2 Model performance evaluation
Mean absolute percentage error(MAPE) is employed to assess the performance of the proposed model. MAPE reflects the average relative deviation of the prediction values from the actual values. Thus the smaller their values, the better the model’s performance. The MAPE of lane $n$ prediction task is the average absolute relative deviation of the prediction performance values from the actual performance values on the lane $n$. The overall MAPE is the average of MAPE on all lanes.
$MAPE_{n}=\frac{1}{M}\sum_{j=1}^{M}\left|\frac{y^{t+1}_{n,j}-{\hat{y}^{t+1}_{n,j}}}{y^{t+1}_{n,j}}\right|\times 100\%$ (15)
$MAPE=\frac{1}{N}\sum_{n=1}^{N}MAPE_{n}$ (16)
where,
$MAPE_{n}$ is the MAPE on lane $n$;
$\hat{y}^{t+1}_{n,j}$ is the predicted pavement performance value of unit $j$ on lane $n$ at time step $t+1$;
$y^{t+1}_{n,j}$ is the actual pavement performance value of unit $j$ on lane $n$ at time step $t+1$;
$MAPE$ is the overall MAPE of all lanes.
@@ 4.3 Model setting
The model parameters can be set as follows. The LSTM-based shared layer includes two layers with a hidden layer of dimension 128. In the multiple task-specific heads, the LSTM contains two layers with a hidden layer of dimension 64, and a fully connected layer has two layers with a hidden layer of dimension 64 and an output layer of dimension 32. The activation function employed in the fully connected layers is the LeakyReLU. The number of task-specific heads is variable, aligning with the number of one-way lanes, and usually ranges from two to four. The fully connected layer in the multiple task-specific output layers is size 32 and output dimension 1. The hyperparameters that underpin the proposed model include the learning rate, batch size, and the number of training epochs that are 0.001, 32, and 200, respectively. The output shape of each layer is shown in Table 3.
@@ 4.4 Model performance
The study encompasses an evaluation of four benchmark models, comprising the Long Short-Term Memory network (LSTM), Gated Recurrent Unit (GRU), eXtreme Gradient Boosting Regressor (XGBoost), and Random Forest (RF). Furthermore, the study involves the establishment of two distinct experimental configurations: the ’lane-specific model’ and the ’mixed model’. The lane-specific model will train a dedicated model for each lane. The ’mix model’ predict pavement performance of all lanes in one unified modeling framework. In all three road scenarios of one-way 2-lane, 3-lane, and 4-lane, the proposed model demonstrates superior performance in terms of MAPE regardless of predicting PCI, PQI, and RQI. The MAPE results of all tests are presented in Table 4.
@@
(1) The proposed model achieves an average of 10.309%, 5.816%, and 4.936% MAPE when predicting PCI, PQI, and RQI on the 2-lane, 3-lane, and 4-lane datasets, respectively. The outcome is related to the quantity of prediction units and tasks. The greater the number of units, the more common and unique pavement decay patterns the model acquires. Furthermore, a greater number of lanes is more conducive to enhancing the process of information sharing.
(2) Compared with the ’lane-specific model’ and ’mix model’, the proposed model has accomplished an average absolute maximum reduction in MAPE of 2.774%, 2.699%, and 3.092% on the 2-lane, 3-lane, and 4-lane datasets, respectively. Correspondingly, in sequence, the average absolute minimum reduction in MAPE is 0.560%, 0.556%, and 1.480% on the datasets mentioned above. The outcomes show the effectiveness of the model in leveraging related tasks to enhance the overall forecasting accuracy in pavement performance analysis. Besides, the MAPE reduction on the 2-lane and 3-lane datasets is closed, while on the 4-lane dataset, the reduction is almost three times that of the reduction on the 2-lane and 3-lane datasets. It demonstrates that increasing the number of tasks does not necessarily increase the magnitude of MAPE reduction.
(3) The model attains an average MAPE of 7.852%, 7.275%, and 5.934%, respectively, for predicting PQI, PCI, and RQI. Furthermore, as depicted in Figure 1, the performance values of different lanes is different, and the dataset exhibits the maximum average PQI difference of 2.834 between different lanes, succeeded by PCI with a difference of 1.950 and RQI with a difference of 0.946. The MAPE demonstrates a specific positive correlation with the performance disparity between lanes. The positive correlation indicates that the similarity between lanes and the resemblance between tasks can mitigate the model’s prediction error.
(4) Figure 4 demonstrates lane-level MAPE results using the proposed model. Moreover, as the number of lanes escalates, the difference in MAPE among lanes narrows. Additionally, the MAPE of lane 1 and lane 2 generally shows a downward trend because the information of the newly added prediction tasks will improve the accuracy of the existing performance prediction tasks.
@@
(5) Figures 5, 6, and 7 show lane-level MAPE based on 2-lane, 3-lane, and 4-lane datasets. In most cases, the MAPE of each lane is the lowest using our model compared with the lane-specific and mix models. In the 2-lane dataset, there are 3 times, accounting for 50%, that the lane-level MAPE is not the lowest. And all are from lane 2. But in 3-lane and 4-lane datasets, the numbers that the lane-level MAPE is not the lowest are 2 and 4, only accounting for 22% and 33%. The result means that the multi-task model is suitable for decreasing the MAPE of each lane.
@@ 4.5 Ablation experiment
An ablation experiment is conducted on the model structure and pavement auxiliary features to assess the distinct contributions of each component on the 4-lane dataset. Regarding the model structure, three different test models were implemented. As for the auxiliary features, each feature was removed individually to assess its impact.
There are test models’ introductions. In test model 1, the LSTM-based shared layer is removed, so the pavement performance time series are sent directly to each task-specific head. The shared layer typically extracts common pavement decay patterns from the input data, offering a universally applicable prediction outcome for all lanes. Furthermore, these layers incorporate a regularization effect that helps prevent over-fitting to any single task. In test model 2, the model eliminates the multiple task-specific heads. Instead, the output from the LSTM-based shared layer is transformed into a 2-dimensional format and then horizontally concatenated with auxiliary features. The presence of task-specific heads, each focusing on the prediction of a specific lane, can enhance prediction performance by concentrating on the lane-specific patterns most indicative of the outcomes for each lane. This specialization improves the accuracy and reliability of the predictions. In test model 3, the horizontal concatenation process is removed. The original model integrates pavement auxiliary features and initial predictions. The former enhances the model’s ability to distinguish different prediction units. It is conducive to capturing the varied impacts on pavement performance deterioration more comprehensively, including factors such as traffic load and pavement material. The latter helps the model to adjust the predictions and get more precise results. It enables the model to detect more subtle and complex patterns. Moreover, feature concatenation provides increased regularization, which helps to mitigate the risk of over-fitting.
The proposed model and test models are used to predict PCI, PQI, and RQI, and the MAPE results are shown in Figure 8(a). The results indicate that all the components contribute positively to enhancing the accuracy of the predictions. The LSTM-based shared layer is of the most significant importance to the proposed model. When this shared layer is removed, the MAPE increases by 1.631%, 1.611%, and 1.806% when predicting PCI, PQI, and RQI, respectively.
To validate the impact of every pavement auxiliary feature on the model’s predictive performance, each auxiliary feature was removed individually, and the MAPE results are shown in Figure 8(b). The results indicate that all the utilized auxiliary features contribute positively to enhancing the accuracy of the predictions. However, there is not one feature that will greatly improve prediction accuracy. The reduction of MAPE ranges from 0.227% to 0.547% when predicting PCI, 0.300% to 0.697% when predicting PQI, and 0.245% to 0.663% when predicting RQI.